OpenStudy (anonymous):

Can someone help me with this... A car traveling at 41 ft/sec decelerates at a constant 4 feet per second per second. How many feet does the car travel before coming to a complete stop?

OpenStudy (anonymous):

i can give it a shot, but no guarantees

OpenStudy (anonymous):

you should post this question in physics group

OpenStudy (anonymous):

acceleration is the second derivative of the position function p, the first derivative is the speed (velocity) v you know that \[a(t)=-4 = v'(t) \] that means \[v(t)=-4t+C\] and we know that the initial speed is 41 so \[v(0)-41\implies v(t)=-4t+41\]

OpenStudy (anonymous):

sorry last line should read \[v(0)=41\implies v(t)=-4t+41\]

OpenStudy (anonymous):

the car will stop when the velocity is 0 so that should be at time \[v(t)=-4t+41=0\] meaning \[t=\frac{41}{4}\]

OpenStudy (anonymous):

now we want the position function, p and we know \[p'(t)=v(t)=-4t+41\] making \[p(t)=-2t^2+41t\]

OpenStudy (anonymous):

correct me if i am making a colossal mistake, i don't know any physics whatsoever

OpenStudy (wasiqss):

this will be goin to b quadritic one

OpenStudy (anonymous):

no , i think now you simply replace t by 41/4 to get the answer

OpenStudy (wasiqss):

v=0, u=-41t

OpenStudy (anonymous):

Uhm, I replaced t in the equation but that answer was wrong... I don't know.

OpenStudy (wasiqss):

sorry dint read full question

OpenStudy (anonymous):

if i am not mistaken now it is a matter of computing \[p(\frac{41}{4})\] so finish. i get 200.125

OpenStudy (anonymous):

actually i get 210.125

OpenStudy (wasiqss):

answer is 10.25

OpenStudy (anonymous):

then i am wrong

OpenStudy (anonymous):

It was 210.125 thanks everyone ^-^

OpenStudy (anonymous):

well, i am not so sure about the 10.25 where did that come from? it seem unlikely to me

OpenStudy (wasiqss):

lolx i was doin programmin so did it wrong lolx